OpenStudy (anonymous):

Write the quadratic function f(x) = 3x^2 - 6x + 9 in vertex form.

OpenStudy (anonymous):

f(x) = 3x² - 6x + 9 f(x) = 3(x² - 2x) + 9 f(x) = 3(x² - 2x + 1) + 9 - [3(1)] f(x) = 3(x - 1)² + 9 - 3 f(x) = 3(x - 1)² + 6

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

Your Welcome:)